OpenStudy (anonymous):

Katrina is fencing her yard, which is shaped as a rhombus. She used (5x + 1) meters of fencing material for one side and (7x - 11) meters for the adjacent side. What is the perimeter of the yard? A.116 meters B.124 meters C.140 meters D.168 meters

OpenStudy (anonymous):

Since a rhombus has 4 equal sides we need to find x first and then the perimeter. 5x+1=7x-11 Then P=2(5x+1)+2(7x-11)

OpenStudy (anonymous):

so its B?

OpenStudy (anonymous):

Correct!!!!